# Learning to Reason for Long-Form Story Generation

### v1

@@ Abstract
Generating high-quality stories spanning thousands of tokens requires competency across a variety of skills, from tracking plot and character arcs to keeping a consistent and engaging style. Due to the difficulty of sourcing labeled datasets and precise quality measurements, most work using large language models (LLMs) for long-form story generation uses combinations of hand-designed prompting techniques to elicit author-like behavior. This is a manual process that is highly dependent on the specific story-generation task. Motivated by the recent success of applying RL with Verifiable Rewards to domains like math and coding, we propose a general story-generation task (Next-Chapter Prediction) and a reward formulation (Verified Rewards via Completion Likelihood Improvement) that allows us to use an unlabeled book dataset as a learning signal for reasoning. We learn to reason over a story’s condensed information and generate a detailed plan for the next chapter. Our reasoning is evaluated via the chapters it helps a story-generator create, and compared against non-trained and supervised finetuning (SFT) baselines. Pairwise human judgments reveal the chapters our learned reasoning produces are preferred across almost all metrics, and the effect is more pronounced in Scifi and Fantasy genres. [^1] We release reproduction & training code: github.com/Alex-Gurung/ReasoningNCP.
@@ 1 Introduction
Long-form story generation is a difficult modeling task that requires synthesizing thousands of tokens of rich and subtle text into coherent and character-aware narratives. High-quality writing balances interesting characters with engaging world-building and satisfying plot arcs (Jarvis, 2014; Kyle, 2016), posing problems of long-term dependencies and accurate Theory-of-Mind modeling. Although Large Language Models (LLMs) have shown recent promise in writing long texts (Yang et al., 2022; 2023; Bai et al., 2024; Xie et al., 2023; Shao et al., 2024a), their stories still struggle on a variety of criteria like originality, plot, character development, and pacing (Chakrabarty et al., 2024; Wang et al., 2023; 2022; Ismayilzada et al., 2024), in addition to more fundamental long-form generation flaws like repetition and quality degradation (Que et al., 2024; Wu et al., 2024).
@@
Outside of long-form story generation, Reinforcement Learning (RL) has been successfully applied to post-training LLMs across various tasks, from optimizing general human preferences to improving performance in specialized domains like code generation and math (Li et al., 2025; Lambert et al., 2024; Kumar et al., 2025). Although some RL work briefly touches on creative writing, these efforts typically focus on short-form tasks and are framed as overall preference alignment (Nguyen et al., 2024; Zhao et al., 2024). In general, RL methods for LLMs have necessitated at least one of the following: (a) high quality datasets labeled with preferences, rewards, or verified reasoning traces or (b) high-quality reward models, often in the form of verifiable reward functions. Large-scale datasets have yielded strong results for broader tasks like human-preference-alignment (Rafailov et al., 2023; Ethayarajh et al., 2024; Cui et al., 2023; Wang et al., 2024a), while recent efforts in math and coding domains have achieved significant success through the use of high-quality rewards (Gehring et al., 2024; DeepSeek-AI et al., 2025; Shao et al., 2024b; Lambert et al., 2024; Kimi Team et al., 2025; Li et al., 2025; Kumar et al., 2025). In particular, there has been a recent surge of interest in enhancing the reasoning capabilities of LLMs using verifiable rewards (Lambert et al., 2024). Instead of relying on a learned reward model, recent work uses deterministic verification functions — usually rule-based methods that assess answers for correctness and formatting.
@@
Neither approach is practical for long-form story generation: story ‘correctness’ is ill-defined and collecting large datasets of labeled story completions is challenging. The multifaceted and nuanced nature of narratives also makes the formulation of singular reward functions difficult. Instead, the ‘quality’ of stories or their continuation is usually assessed through pairwise comparisons across multiple metrics (Yang et al., 2022; Huot et al., 2024; Yang et al., 2023; Xie & Riedl, 2024). Most current work on story generation has therefore avoided RL, and instead leaned into hand-designed systems that mimic different parts of the human writing process like drafting, editing, and planning. (Fan et al., 2018; Zhou et al., 2023; Wang et al., 2023; Xie & Riedl, 2024; Wang et al., 2024b). Huot et al. (2024) train LLMs to model human story preferences across metrics like creativity and language use, but they do not train story generation models using these evaluators.
@@
Rather than predicting and evaluating entire book-length generations, we propose Next-Chapter Prediction (NCP) as a more tractable and informative task. Inspired by the human book-writing process (Kyle, 2016; Jarvis, 2014) that uses both a high-level sketch of the story and more fine-grained information about characters and plot progression, we model story generation as predicting the next chapter given a similar collection of story information. Finetuning directly on this task quickly leads to overfitting and fails to capture the underlying reasoning behind the story-writing process. Instead, we propose a novel reasoning paradigm akin to Reinforcement Learning from Verifiable Rewards (RLVR; Lambert et al. 2024) that allows RL training of reasoning traces for story generation.
@@
We introduce Verifiable Rewards via Completion Likelihood Improvement (VR-CLI), a reward modeling paradigm designed to learn reasoning traces that enhance a generator’s ability to reproduce a given dataset (see Figure 1). Our key assumption is that increasing the generator’s likelihood of predicting the next chapter will in turn improve the quality of its generations. Accordingly, our reward is defined with respect to the improvement in predicting a gold next chapter, which can be naturally expressed in terms of per-token perplexity. We use VR-CLI to train Qwen 2.5 models (Qwen et al., 2024) to produce traces that cite and reason about the given story information before constructing a detailed plan, which is then fed as context to the base model for generating the chapter continuation. Our results show that these models outperform reasoning and non-reasoning baselines and models trained via SFT on next-chapter prediction. Our contributions are as follows:
@@ 2 Related Work
Story generation has long been studied as a completion task (Mostafazadeh et al., 2016; Fan et al., 2018), but only with recent advances in long-context modeling have language models begun producing stories of considerable length. A large body of previous work has shown that incorporating condensed story information, such as plot outlines (Fan et al., 2018; Zhou et al., 2023; Wang et al., 2023; Xie & Riedl, 2024; Wen et al., 2023; Yoo & Cheong, 2024), as well as setting and character details (Yang et al., 2022; 2023), significantly improves generation quality. Further improvements have been achieved by explicitly breaking down the writing process into sub-tasks handled by different ‘agents’ or models (Huot et al., 2024; Peng et al., 2022). However, these approaches largely rely on hand-crafted prompting techniques to generate and refine both the condensed story information and the story itself. Progress in long-form story generation has been hindered by challenges in evaluation which is inherently subjective, nuanced, and time-consuming. Human evaluation has become the de facto standard for assessing machine-generated stories, typically through pairwise judgments across key dimensions such as coherence, plot development, creativity, and characterization (Yang et al., 2022; 2023; Huot et al., 2024; Xie & Riedl, 2024; Chhun et al., 2022).
@@
Finetuning LLMs with Reinforcement Learning (RL) has become an increasingly popular method for improving performance on a variety of tasks. Traditionally, RL has been used to align model generations with human preferences (RLHF; Grattafiori et al. 2024; Achiam et al. 2023; Ziegler et al. 2019; Ouyang et al. 2022; Stiennon et al. 2020; Christiano et al. 2017). Building on these online methods, recent work has introduced algorithms that can also learn policies from static preference datasets (Rafailov et al., 2023; Ethayarajh et al., 2024).
RL techniques have further led to significant advancements in verifiable and executable domains such as math, science, and coding (Gehring et al., 2024; Simonds & Yoshiyama, 2025; Li et al., 2025; Kumar et al., 2025). Several recent large-model releases have also emphasized the use of RL to improve performance on math and programming benchmarks (DeepSeek-AI et al., 2025; Kimi Team et al., 2025). Many of these approaches build upon the Reinforcement Learning from Verifiable Rewards (RLVR) paradigm introduced in Lambert et al. (2024) which trains LLMs to produce useful reasoning traces for tasks with easily verified answers via a binary or scaled reward.
We are not aware of previous work that learns reasoning traces for creative long-form generation, possibly due to the challenge of defining rewards with objective correctness criteria. However most similar in approach, Hu et al. (2023) use LLM-based GFLowNets (Bengio et al., 2021) for sentence continuation and 5-sentence-story infilling using the ROCStories dataset (Mostafazadeh et al., 2016). By representing many LLM tasks as latent variable modeling problems they connect RL approaches to the field of probabilistic inference.
@@ 3 Next-Chapter Prediction
[…]
@@
At a chapter index $i$ with Story-Information $SI_{i}$, we denote the predicted next chapter as $\hat{c}_{i+1}$. When the next-chapter is given, not predicted, we denote it $c_{i+1}$. We denote the story-generator model that predicts the next chapter $\pi^{\mathcal{G}}_{\theta}$. The default setup for this task is to directly generate the next chapter from the story-information, using the story-generator:
@@
Our claim in this work is that reasoning traces can lead to better next-chapter prediction; specifically, we use a reasoning model $\pi^{\mathcal{R}}_{\theta}$ to predict a reasoning trace about the given story information, ending with a more detailed plan for the next chapter. We denote this predicted plan $\hat{p}$. In reasoning model variants we generate the next chapter using a story-generator model conditioned on both the story-information and the plan:
@@
$\hat{c}_{i+1}\leftarrow\pi^{\mathcal{G}}_{\theta}(SI_{i},\hat{p})$ (3)
We evaluate the effectiveness of story-generation and reasoning models via the quality of the next chapter, judged both on its own merits and its ability to fit into the broader story context. More details are provided in Section 7.2, and example reasoning traces are in Appendix G.
@@ 4 Dataset Curation
We collect a dataset of 30 books published in or after 2024 (to mitigate information leakage). [^2] Our work primarily uses two models: Llama 3.3 70B and Qwen 2.5 3B/7B-1M Instruct. The Llama model has a knowledge cutoff of December 2023, and while the Qwen models were released in 2024 we find little evidence of data contamination. Details on book selection are in Appendix A.1. Our books range from 67k to 214k tokens, with a mean of 139k. Rather than raw story text we operate on higher-level plot and character representations: prior research has shown these to be useful for story tasks (Pham et al., 2025; Gurung & Lapata, 2024). We follow previous work in sourcing gold-standard chapter summaries from SuperSummary [^3] https://supersummary.com/ (Xu et al., 2024; Yuan et al., 2024; Ou & Lapata, 2025). We explain below how we use these in building the story-information $SI_{i}$.
@@ 4.1 Story Information
We denote the following story-information $SI_{i}$, defined at a given story and chapter index $i$.
High-level Story Sketch This can be thought of as an author’s sketch of how the story should unfold. As we do not have these sketches available, we mimic them by summarizing all of a books individual chapter summaries together into a high-level plan (see Table 1). We use Llama 3.3 70B (Grattafiori et al., 2024) to generate these summaries.
@@
Character Sheets Prior work has shown character information is useful for downstream story tasks (Gurung & Lapata, 2024; Huot et al., 2024), so we incorporate character information into $SI_{i}$ by adapting the CHIRON character sheets for longer stories (Gurung & Lapata, 2024). These sheets contain four broad categories of character information and are automatically generated from story text before being filtered for accuracy using an entailment module. We create individual character sheets for the three main characters of each story, defined at each chapter index $i$ based on the previous chapters $\leq i$. We use the Llama 3.3 70B model for both generation and entailment modules and add a summarization step to consolidate the sheets. More details are in Appendix A.3.
Next Chapter Synopsis For a given chapter index $i$, we call the summary of the next chapter a ‘synopsis’. This contains a rough idea of what should happen in the chapter, which informs the path the story generations take. On average, the synopsis is 7.4% of the size of the next chapter (in tokens). We hypothesize that by adding detail to the synopsis via reasoning (see Table 1), we can more effectively guide the generation for the chapter.
@@ 4.2 Dataset Statistics
[…]
@@ 5 Verifiable Rewards via Completion Likelihood Improvement (VR-CLI)
Taking inspiration from RLVR (Lambert et al., 2024) and the latent-variable modeling framing of Hu et al. (2023), we introduce Verifiable Rewards via Completion Likelihood Improvement (VR-CLI). VR-CLI is a reward modeling paradigm that learns reasoning traces that help a generator reproduce a given dataset without the need for labeled data or defined external rewards. This formulation relies on a key assumption: improving the generator’s likelihood of producing the provided dataset will in turn improve the quality of its generations. We verify this assumption for NCP in Section 8. We believe VR-CLI is a promising reward paradigm for other tasks where completions are not easily verified.
Let $D$ represent a dataset of pairs $(x,y)$, where $x$ is the input prompt, and $y$ is the gold completion. A reasoning model generates a reasoning trace $t$ which culminates in a final answer $a$. We define a generator model, $\pi^{\mathcal{G}}$, that is not trained but instead used to get the likelihood of generating $y$. For our NCP task, $x$ is the story-information, $y$ is the gold next chapter, $a$ is the detailed plan $\hat{p}$ and our generator $\pi^{\mathcal{G}}$ is the story generator model.
@@
Positive values imply the perplexity with reasoning is lower than the perplexity without, and the reasoning improved the likelihood of generating the gold continuation. Negative values imply the opposite, that the reasoning worsened the likelihood. Note that $PPL_{\pi^{\mathcal{G}}}(y|x)$ does not depend on the policy model or answers, so it can be pre-computed before training.
There are two ways to define our reward based on the improvement $I$: (1) as a piecewise function, like RLVR, subject to the magnitude of the improvement and custom thresholds ($\{\omega_{0},\omega_{1},...\omega_{n}\}$) or (2) as a raw value, bounded by 0 when necessary. $R(x,y,a)=\begin{cases}\alpha&I(x,y,a)\leq\omega_{0}\\ \beta&\omega_{0}<I(x,y,a)\leq\omega_{1}\\ \gamma&\omega_{1}<I(x,y,a)\\ \end{cases}$ (5) $R(x,y,a)=\max[0,I(x,y,a)]$ (6)

### v2

@@ Abstract
Generating high-quality stories spanning thousands of tokens requires competency across a variety of skills, from tracking plot and character arcs to keeping a consistent and engaging style. Due to the difficulty of sourcing labeled datasets and precise quality measurements, most work using large language models (LLMs) for long-form story generation resorts to combinations of hand-designed prompting techniques to elicit author-like behavior. This is a manual process that is highly dependent on the specific story-generation task. Motivated by the recent success of applying RL with Verifiable Rewards to domains like math and coding, we propose a general story-generation task (Next-Chapter Prediction) and a reward formulation (Verifiable Rewards via Completion Likelihood Improvement) that allows us to use an unlabeled book dataset as a learning signal for reasoning. We learn to reason over a story’s condensed information and generate a detailed plan for the next chapter. Our reasoning is evaluated via the chapters it helps a story generator create, and compared against non-trained and supervised fine-tuning (SFT) baselines. Pairwise human judgments reveal the chapters our learned reasoning produces are preferred across almost all metrics, and the effect is more pronounced in Sci-Fi and Fantasy genres. [^1] We release reproduction and training code at github.com/Alex-Gurung/ReasoningNCP.
@@ 1 Introduction
Long-form story generation is a difficult modeling task that requires synthesizing thousands of tokens of rich and subtle text into coherent and character-aware narratives. High-quality writing balances interesting characters with engaging world-building and satisfying plot arcs (Jarvis, 2014; Kyle, 2016), posing problems of long-term dependencies and accurate Theory-of-Mind modeling. Although Large Language Models (LLMs) have shown recent promise in writing long texts (Yang et al., 2022; 2023; Bai et al., 2025; Xie et al., 2023; Shao et al., 2024a), their stories still struggle on a variety of criteria like originality, plot, character development, and pacing (Chakrabarty et al., 2024; Wang et al., 2023; 2022; Ismayilzada et al., 2024), in addition to more fundamental long-form generation flaws like repetition and quality degradation (Que et al., 2024; Wu et al., 2024).
@@
Outside of long-form story generation, Reinforcement Learning (RL) has been successfully applied to post-training LLMs across various tasks, from optimizing general human preferences to improving performance in specialized domains like code generation and math (Li et al., 2025; Lambert et al., 2024; Kumar et al., 2025). Although some RL work briefly touches on creative writing, these efforts typically focus on short-form tasks and are framed as an overall preference alignment (Nguyen et al., 2024; Zhao et al., 2024). In general, RL methods for LLMs have necessitated at least one of the following: (a) high quality datasets labeled with preferences, rewards, or verified reasoning traces or (b) high-quality reward models, often in the form of verifiable reward functions. Large-scale datasets have yielded strong results for broader tasks like human-preference-alignment (Rafailov et al., 2023; Ethayarajh et al., 2024; Cui et al., 2024; Wang et al., 2025a), while recent efforts in the math and coding domains have achieved significant success through the use of high-quality rewards (Gehring et al., 2025; DeepSeek-AI et al., 2025; Shao et al., 2024b; Lambert et al., 2024; Kimi Team et al., 2025; Li et al., 2025; Kumar et al., 2025). In particular, there has been a recent surge of interest in enhancing the reasoning capabilities of LLMs using verifiable rewards (Lambert et al., 2024). Instead of relying on a learned reward model, recent work uses deterministic verification functions — usually rule-based methods that assess answers for correctness and formatting.
@@
Neither approach is practical for long-form story generation: story ‘correctness’ is ill-defined, and collecting large datasets of labeled story completions is challenging. The multifaceted and nuanced nature of narratives also makes the formulation of singular reward functions difficult. Instead, the ‘quality’ of stories or their continuation is usually assessed through pairwise comparisons across multiple metrics (Yang et al., 2022; Huot et al., 2025; Yang et al., 2023; Xie & Riedl, 2024). Most current work on story generation has therefore avoided RL, and instead leaned into hand-designed systems that mimic different parts of the human writing process, like drafting, editing, and planning. (Fan et al., 2018; Zhou et al., 2023; Wang et al., 2023; Xie & Riedl, 2024; Wang et al., 2025b). Huot et al. (2025) train LLMs to model human story preferences across metrics like creativity and language use, but they do not train story generation models using these evaluators.
@@
Rather than predicting and evaluating entire book-length generations, we propose Next-Chapter Prediction (NCP) as a more tractable and informative task. Inspired by the human book-writing process (Kyle, 2016; Jarvis, 2014) that uses both a high-level sketch of the story and more fine-grained information about characters and plot progression, we model story generation as predicting the next chapter given a similar collection of story information. Fine-tuning directly on this task, quickly leads to overfitting and fails to capture the underlying reasoning behind the story-writing process. Instead, we propose a novel reasoning paradigm akin to Reinforcement Learning from Verifiable Rewards (RLVR; Lambert et al. 2024) that allows RL training of reasoning traces for story generation.
@@
We introduce Verifiable Rewards via Completion Likelihood Improvement (VR-CLI), a reward modeling paradigm designed to learn reasoning traces that enhance a generator’s ability to reproduce a given dataset (see Figure 1). Our key assumption is that increasing the generator’s likelihood of predicting the next chapter will, in turn, improve the quality of its generations. Accordingly, our reward is defined with respect to the improvement in predicting a gold next chapter, which can be naturally expressed in terms of per-token perplexity. We use VR-CLI to train Qwen 2.5 models (Qwen et al., 2024) to produce traces that cite and reason about the given story information before constructing a detailed plan, which is then fed as context to the base model for generating the chapter continuation. Our results show that these models outperform reasoning and non-reasoning baselines and models trained via SFT on next-chapter prediction. Our contributions are as follows:
@@ 2 Related Work
Story generation has long been studied as a completion task (Mostafazadeh et al., 2016; Fan et al., 2018), but only with recent advances in long-context modeling have language models begun producing stories of considerable length. A large body of previous work has shown that incorporating condensed story information, such as plot outlines (Fan et al., 2018; Zhou et al., 2023; Wang et al., 2023; Xie & Riedl, 2024; Wen et al., 2023; Yoo & Cheong, 2024), as well as setting and character details (Yang et al., 2022; 2023), significantly improves generation quality. Further improvements have been achieved by explicitly breaking down the writing process into sub-tasks handled by different ‘agents’ or models (Huot et al., 2025; Peng et al., 2022). However, these approaches largely rely on hand-crafted prompting techniques to generate and refine both the condensed story information and the story itself. Progress in long-form story generation has been hindered by challenges in evaluation, which is inherently subjective, nuanced, and time-consuming. Human evaluation has become the de facto standard for assessing machine-generated stories, typically through pairwise judgments across key dimensions such as coherence, plot development, creativity, and characterization (Yang et al., 2022; 2023; Huot et al., 2025; Xie & Riedl, 2024; Chhun et al., 2022).
@@
Fine-tuning LLMs with Reinforcement Learning (RL) has become an increasingly popular method for improving performance on a variety of tasks. Traditionally, RL has been used to align model generations with human preferences (RLHF; Grattafiori et al. 2024; Achiam et al. 2023; Ziegler et al. 2019; Ouyang et al. 2022; Stiennon et al. 2020; Christiano et al. 2017). Building on these online methods, recent work has introduced algorithms that can also learn policies from static preference datasets (Rafailov et al., 2023; Ethayarajh et al., 2024).
RL techniques have further led to significant advancements in verifiable and executable domains such as math, science, and coding (Gehring et al., 2025; Simonds & Yoshiyama, 2025; Li et al., 2025; Kumar et al., 2025). Several recent large-model releases have also emphasized the use of RL to improve performance on math and programming benchmarks (DeepSeek-AI et al., 2025; Kimi Team et al., 2025). Many of these approaches build upon the Reinforcement Learning from Verifiable Rewards (RLVR) paradigm introduced in Lambert et al. (2024), which trains LLMs to produce useful reasoning traces for tasks with easily verified answers via a binary or scaled reward.
We are not aware of previous work that learns reasoning traces for creative long-form generation, possibly due to the challenge of defining rewards with objective correctness criteria. However most similar in approach, Hu et al. (2024a) use LLM-based GFLowNets (Bengio et al., 2021) for sentence continuation and 5-sentence-story infilling using the ROCStories dataset (Mostafazadeh et al., 2016). By representing many LLM tasks as latent variable modeling problems, they connect RL approaches to the field of probabilistic inference.
Concurrent work has also begun exploring the idea of using a model’s likelihood of the solution to optimize reasoning. JEPO (Tang et al., 2025), VeriFree (Zhou et al., 2025), LATRO (Chen et al., 2024) and INTUITOR (Zhao et al., 2025) all propose similar policy optimizers based on this latent-variable perspective, and largely apply their methods to math tasks. Our VR-CLI formulation differs in its reward shaping, lack of SFT-objective, and its application to long-form creative writing.
@@ 3 Next-Chapter Prediction
Next Chapter And when a voice came through the wires, it wasn’t a member of the staff, but Peter Van Laar himself—to whom Carl nodded each time they crossed paths at the Preserve, but to whom he had actually spoken maybe twice in his life.
@@
At a chapter index $i$ with Story-Information $SI_{i}$, we denote the predicted next chapter as $\hat{c}_{i+1}$. When the next chapter is given, not predicted, we denote it $c_{i+1}$. We denote the story-generator model that predicts the next chapter $\pi^{\mathcal{G}}_{\theta}$. The default setup for this task is to directly generate the next chapter from the story information, using the story generator:
@@
Our claim in this work is that reasoning traces can lead to better next-chapter prediction; specifically, we use a reasoning model $\pi^{\mathcal{R}}_{\theta}$ to predict a reasoning trace about the given story information, ending with a more detailed plan for the next chapter. We denote this predicted plan $\hat{p}$. In reasoning model variants, we generate the next chapter using a story-generator model conditioned on both the story information and the plan:
@@
[…]
We evaluate the effectiveness of story-generation and reasoning models via the quality of the next chapter, judged both on its merit and its ability to fit into the broader story context. Note that our proposed reasoning method optimizes the detailed plans, which are then used to improve the generated chapters. More details are provided in Section 7.2, and example reasoning traces are in Appendix G.
@@ 4 Dataset Curation
We collect a dataset of 30 books published in or after 2024 (to mitigate information leakage). [^2] Our work primarily uses two models: Llama 3.3 70B and Qwen 2.5 3B/7B-1M Instruct. The Llama model has a knowledge cutoff of December 2023, and while the Qwen models were released in 2024, we find little evidence of data contamination. Details on book selection are in Appendix A.1. Our books range from 67k to 214k tokens, with a mean of 139k. Rather than raw story text we operate on higher-level plot and character representations as prior research has shown these to be useful for story tasks (Pham et al., 2025; Gurung & Lapata, 2024). We follow previous work in sourcing gold-standard chapter summaries from SuperSummary [^3] https://supersummary.com/ (Xu et al., 2024; Yuan et al., 2024; Ou & Lapata, 2025). We explain below how we use these in building the story-information $SI_{i}$.
@@ 4.1 Story Information
We denote the following story information $SI_{i}$, defined at a given story and chapter index $i$.
High-level Story Sketch This can be thought of as an author’s sketch of how the story should unfold. As we do not have these sketches available, we mimic them by summarizing all of a book’s individual chapter summaries together into a high-level plan (see Table 1). We use Llama 3.3 70B (Grattafiori et al., 2024) to generate these summaries.
@@
Character Sheets Prior work has shown that that character information is useful for downstream story tasks (Gurung & Lapata, 2024; Huot et al., 2025), so we incorporate character information into $SI_{i}$ by adapting the CHIRON character sheets for longer stories (Gurung & Lapata, 2024). These sheets contain four broad categories of character information and are automatically generated from story text before being filtered for accuracy using an entailment module. We create individual character sheets for the three main characters of each story, defined at each chapter index $i$ based on the previous chapters $\leq i$. We use the Llama 3.3 70B model for both generation and entailment modules and add a summarization step to consolidate the sheets. More details are in Appendix A.3.
Next Chapter Synopsis For a given chapter index $i$, we call the summary of the next chapter a ‘synopsis’. This contains a rough idea of what should happen in the chapter, which informs the path the story generations take. On average, the synopsis is 7.4% of the size of the next chapter (in tokens). We hypothesize that by adding detail to the synopsis via reasoning (see Table 1), we can more effectively guide the generation of the chapter.
@@ 4.2 Dataset Statistics
The final dataset has 30 books, split 22-4-4 into training, validation, and testing. We split by book to evaluate our method’s generalization capabilities across books and authors, as training may learn book or author-specific style or plot information. Further splitting into chapters and lightly filtering for chapter length ($200\leq\text{\# words}\leq 5,000$) and chapter location (2<i<|S|22<i<|S|-2) gives us 1,004 training datapoints, 162 for validation, and 181 for testing. More details are provided in Appendix A.4 and Tables 6, and 7. Example sentences from each element in our dataset are in Table 1.
@@ 5 Verifiable Rewards via Completion Likelihood Improvement (VR-CLI)
Taking inspiration from RLVR (Lambert et al., 2024) and the latent-variable modeling framing of Hu et al. (2024a), we introduce Verifiable Rewards via Completion Likelihood Improvement (VR-CLI). VR-CLI is a reward modeling paradigm that learns reasoning traces that help a generator reproduce a given dataset without the need for labeled data or defined external rewards. This formulation relies on a key assumption: improving the generator’s likelihood of producing the provided dataset will, in turn, improve the quality of its generations. We verify this assumption for NCP in Section 8. We believe VR-CLI is a promising reward paradigm for other tasks where completions are not easily verified.
Let $D$ represent a dataset of pairs $(x,y)$, where $x$ is the input prompt, and $y$ is the gold completion. A reasoning model generates a reasoning trace $t$ which culminates in a final answer $a$. We define a generator model, $\pi^{\mathcal{G}}$, that is not trained but instead used to get the likelihood of generating $y$. For our NCP task, $x$ is the story information, $y$ is the gold next chapter, $a$ is the detailed plan $\hat{p}$, and our generator $\pi^{\mathcal{G}}$ is the story generator model.
@@
Positive values imply the perplexity with reasoning is lower than the perplexity without, and the reasoning improved the likelihood of generating the gold continuation. Negative values imply the opposite, that the reasoning worsened the likelihood. Note that $PPL_{\pi^{\mathcal{G}}}(y|x)$ does not depend on the policy model or answers, so it can be pre-computed before training. Although we use perplexity as our likelihood measure in this work, a very similar formulation could be constructed with average log-likelihood instead. We found both methods to produce similar results; more details are presented in Appendix A.8.
[…]
@@
Additionally, one could forgo the baseline and use the chosen likelihood-measure directly (e.g. $I_{\pi^{\mathcal{G}}}(x,y,a)=-PPL(y|x,a)$). Notably, when using GRPO’s advantage estimation and a simple reward equal to improvement ($R(x,y,a)=I(x,y,a)$), advantages are equivalent with and without the baseline (Shao et al., 2024b). However, we hypothesize that including a baseline provides a more interpretable and useful reward estimate, especially when using a more complicated reward definition. For example, across a group where all samples produce worse-than-baseline perplexities, it may be useful to set their rewards uniformly to zero (and not increase the likelihood of potentially damaging traces). We also test this in ablations in Appendix A.8 and found comparable results, and believe both formulations are worth exploring in future tasks. As all of these formulations retain the goal of increasing the gold-completion’s likelihood via the reward, we view them as variants of the VR-CLI paradigm.